OpenStudy (ccswims):

WILL MEDAL Prove the Converse of the Pythagorean Theorem using similar triangles. The Converse of the Pythagorean Theorem states that when the sum of the squares of the lengths of the legs of the triangle equals the squared length of the hypotenuse, the triangle is a right triangle. Be sure to create and name the appropriate geometric figures.

OpenStudy (ccswims):

@Hero can you help me?

OpenStudy (anonymous):

@ccswims I can prove this one for you but I am sleepy so give me a shake and hug and I will help you

OpenStudy (anonymous):

..........

OpenStudy (ccswims):

? wait what do you mean?

OpenStudy (anonymous):

Ok just type "Robert I need your help don't fokiing sleep and help me"

OpenStudy (ccswims):

Robert, I need your help, drink an energy drink if you have to but dont fall asleep on me

OpenStudy (anonymous):

Ok no problem

OpenStudy (anonymous):

First off in proving this theory you need to construct squares above the triangle.

OpenStudy (ccswims):

ok, how?

OpenStudy (anonymous):

Ok c^2=4(ab/2)+(a-b)^2=a^2+b^2

OpenStudy (anonymous):